# Mingyu Liu1    Jiuhe Shu1∗    Hui Chen1    Zeju Li1    Canyu Zhao1 Jiange Yang2    Shenyuan Gao3    Hao Chen1    Chunhua Shen1

### v1

@@ Abstract
A fundamental challenge in embodied intelligence is developing expressive and compact state representations for efficient world modeling and decision making. However, existing methods often fail to achieve this balance of compactness and expressivity, yielding representations that are either overly redundant or lacking in task-critical information. We propose an unsupervised approach that learns a highly compressed two-token state representation using a lightweight encoder and a pre-trained Diffusion Transformer (DiT) decoder, capitalizing on its strong generative prior. Our representation is efficient, interpretable, and integrates seamlessly into existing VLA-based models, improving performance by 14.3% on LIBERO and 30% in real-world task success rate with minimal inference overhead. More importantly, we find that the difference between these tokens, obtained via latent interpolation, naturally represents the motion, which can be further decoded into executable robot actions. This emergent capability reveals that our representation captures dynamics without explicit motion supervision. We name our method StaMo for its ability to learn generalizable robotic Motion from compact State representation, which is encoded from static images, challenging the prevalent dependence to learning robotic motions with complex temporal modeling and video data. Our learned representations also enhance policy co-training, outperforming prior methods by 10.4% with improved interpretability. Moreover, our approach scales effectively across diverse data sources, including real-world robot data, simulation, and human egocentric video. Code available at aim-uofa.github.io/StaMo .
@@ 1 Introduction
Learning reusable and generalizable representations is a cornerstone of intelligent robotics systems. While visual representations in VLAs retain rich perceptual details for multimodal fusion, state representations for world modeling and intermediate reasoning serve a different purpose: they are design to enable efficient future prediction and bridging visual planning with action execution. This role demands two key properties: First, it must be highly compact. Unlike rich visual embeddings used for perception, a state representation’s proximity to action generation demands a focus on actionable information to ensure efficient future prediction. Second, it must be expressive despite its compactness. Common low-dimensional approaches, such as trajectories (Wen et al., 2023) or flow fields (Gao et al., 2024; Xu et al., ), often fail on this front; they capture basic motion but lack the semantic richness to encode goal states, interaction dynamics, or structured spatial relationships.
Building on these principles, we propose a compact state representation learned using a Diffusion Autoencoder fine-tuned on robotics data. Inspired by prior work (Zhao et al., 2024) that applied Diffusion Autoencoders to facilitate long video generation in VLMs, our method compresses an image into a highly compact latent state (as few as 2 tokens of 1024 dimensions). To ensure the representation is rich, we initialize its decoder from a powerful DiT model pretrained on internet data (Esser et al., 2024), reasoning that the ability to reconstruct pixels accurately requires an implicit understanding of key state information like robot poses and object interactions. This representation integrates seamlessly into existing VLM architectures, extensive experiments show that our method significantly enhances model performance with negligible overhead to inference speed, achieving improvements of + 14.3% in LIBERO and + 30% in real-world deployments.
Interestingly, from this compact state representation, we make a important discovery: robot motion (which is also known as latent action) naturally emerges from the state representation space. By simply performing a linear interpolation between the latent encodings of start and goal observations, we can generate smooth, plausible, and dynamically consistent motion trajectories. This observation provides an alternative to the dominant paradigm of learning latent actions from video data (Ye et al., 2024a; Bu et al., 2025b; Yang et al., 2025; Chen et al., 2024a). While using video seems intuitive as action is inherently temporal, it presents significant challenges: it demands complex, computationally expensive temporal models and often produces ambiguous, coarse-grained actions. This is because high-variance motion within video clips causes models to learn averaged-out representations, making per-frame learning both inefficient and representationally flawed.
@@ 2 Related Works
2.1 Robot-centric Representation Learning
Previous robotics research on visual representations has often faced a fundamental trade-off: methods excelling at motion representation, like latent actions (Bu et al., 2025b; Ye et al., 2024a; Cheang et al., 2024b; Bjorck et al., 2025; Gao et al., 2025), flow (Gao et al., 2024; Xu et al., 2024), or trajectories (Wen et al., 2023), typically lack a rich, compact state representation. Conversely, approaches that encode detailed state information from raw images (Wang et al., 2025; Bharadhwaj et al., 2024) or dense features (Zhang et al., 2025; Li et al., 2025a; Nair et al., 2022; Xiao et al., 2022; Majumdar et al., 2023; Radosavovic et al., 2023) are often too high-dimensional and redundant to effectively represent motion via simple differences. StaMo, as shown in Figure 2, overcomes this dichotomy by achieving a unique balance. It learns to compress and encode a robot’s visual state into a highly compact token space that is both expressive enough for complex tasks and minimal enough that motion can be elegantly and powerfully represented as the difference between two states. This unified approach not only enables efficient world modeling but also demonstrates superior generalization and scalability, providing a robust foundation for future robotic systems.
2.2 World Modeling in Vision-Language-Ation Models
Benefiting from powerful vision-language models (VLMs) (Bai et al., 2025; Chen et al., 2024b; Beyer et al., 2024) pre-trained on large-scale internet data, vision-language-action (VLA) models have demonstrated significant potential in generalizing across manipulation tasks. Typically, VLAs (Brohan et al., 2022; Zitkovich et al., 2023; Cheang et al., 2024a; Ha et al., 2023; Kim et al., 2024; Black et al., ; Wen et al., 2025; Team et al., 2024) map visual information and language instructions into the robot’s action space through end-to-end training. A natural extension is to endow the model with world modeling capabilities. Inspired by recent unified generative and comprehension models, a promising direction is to enable large models to reason with images—predicting future visual states while inferring actions or language, thereby creating a mutually reinforcing process. Several works (Wang et al., 2025; Zhang et al., 2025; Cen et al., 2025; Li et al., 2025b) have explored this idea. However, these approaches either require decoding full images during inference or rely on overly redundant state representations, limiting their generalization to novel scenes. In contrast, our method employs a more compact representation that improves model performance with minimal impact on inference speed. Furthermore, our representation exhibits strong generalization ability, enabling adaptation to unseen scenes without requiring further fine-tuning of the encoder.
@@ 2.3 Latent Action Learning
Robot learning scalability is constrained by scarce and heterogeneous robotic data, prompting the use of large-scale, action-free internet videos to instill foundational physical and operational knowledge for improved generalization and data efficiency (McCarthy et al., 2024). Numerous studies have explored learning discrete (Ye et al., 2024a; Chen et al., 2024a; Bruce et al., 2024; Bu et al., 2025b; Schmidt & Jiang, 2023; Bu et al., 2025a) or continuous (Yang et al., 2025; Gao et al., 2025) latent actions from unlabeled data and have demonstrated their effectiveness across a wide range of downstream tasks. However, most of these approaches rely on carefully designed model architectures and depend on extracting frames from continuous videos, which introduces limitations such as sensitivity to frame sampling intervals, potential temporal biases, and poorly interpretable, blurry action representations. In contrast, our approach reveals that actionable representations can inherently emerge from large-scale natural image collections—challenging the conventional belief that latent actions must be learned from video sequences. We further demonstrate that such image-emergent actions exhibit superior interpretability and stronger generalization capabilities.
@@ 3 Method
3.1 Compress Images into Compact Representations
@@
To this end, as illustrated in Figure 1, the encoder of our Diffusion Autoencoder, denoted as $\mathcal{E}$, consists of a pretrained DINOv2 model (Oquab et al., 2023) serving as a frozen feature extractor, which is followed by a transformer-based compressor, to map the input observations into a sequence of compact token representations. The decoder $\mathcal{D}$, is a Diffusion Transformer (DiT) (Peebles & Xie, 2023) conditioned on these tokens to reconstruct the original image observations. Our implementation is built upon the Stable Diffusion 3, where only the transformer-based compressor and the DiT decoder are trained. The DINOv2 encoder remains frozen throughout the training process. We employ the same Flow Matching (Albergo & Vanden-Eijnden, 2022; Liu et al., 2022; Lipman et al., 2022) objective used in base model training to optimize the model, which minimizes the loss between the velocity predicted by $\mathcal{D}$ and the target velocity $u$:
@@
where $\mathbf{x}_{0}$ is the input image observation and $\tau$ is the VAE encoder of pretrained diffusion model to produce latent $\mathbf{z}_{0}$. $\mathbf{z}_{t}=(1-\sigma_{t})\mathbf{z}_{0}+\sigma_{t}\epsilon$ is the linear interpolation between pure noise $\epsilon$ and $\mathbf{z}_{0}$. After training, The trained encoder $\mathcal{E}$ is able to yield the state representation $s$. Please refer to Appendix A for more details about architecture and training process.
3.2 Compact Representations for Both State and Motion
@@ 3.3 StaMo for Efficient World Modeling
$L_{total}=\lambda_{action}L_{action}+\lambda_{future}(L_{mse}(s_{pred},s_{gt})+L_{1}(s_{pred},s_{gt})),$ (2)
@@ 3.4 StaMo for Latent Motions Co-Training
We tested this co-training framework with both in-domain robot videos and out-of-domain human demonstration videos to assess the generalization of our motion representation. As shown in Section 4.3, this method leads to substantial performance gains compared with previous method. The results confirm that the motions emerging from StaMo serve as an effective and robust proxy for true actions, providing a scalable pathway to enhance policy learning by leveraging the wealth of unlabeled visual data readily available.
@@ 4.1 Qualitative Analysis of STAMO
In this subsection, we present a qualitative visualization and analysis of the StaMo results. StaMo demonstrates the ability to perform high-quality reconstructions of robotic manipulation images with strong generalization capabilities. It achieves excellent results on both in-domain and out-of-domain data. Quantitative reconstruction results are shown in Table 1, and some qualitative results are presented in Figure 3.
@@ 4.2 World Modeling Results
Table 2: Inference frequency.
Method Frequency (Hz)
UniVLA 2.65
OpenVLA 4.16
@@
In this section, we evaluate the application of StaMo for efficient world modeling. A key advantage of our approach is its ability to integrate seamlessly into existing VLM-based frameworks. We demonstrate this by conducting experiments on two strong baselines: OpenVLA and its successor, OpenVLA-OFT. We specifically chose OpenVLA-OFT because it introduces architectural improvements like parallel decoding and action chunking, which result in a significantly longer action horizon. This requires the model to predict a state representation at a more distant future step, making it a more challenging and relevant benchmark for world modeling. We reproduced the OpenVLA-OFT model and removed the wrist camera input.
Both the StaMo state and motion representations can serve as predictive targets for a world model, and their performance is detailed in Table 4. Our findings indicate that the optimal choice of representation depends on the policy’s action horizon. For the standard OpenVLA, which predicts actions over a short, single-step interval, the motion representation yields better performance. This is because the latent motion is conceptually analogous to the delta end-effector (EE) pose used in frameworks like LIBERO, providing a direct and incremental target. Conversely, for OpenVLA-OFT, which plans over a longer execution step, the state representation is recommended, as it acts as a stable goal-conditioning signal. Importantly, as we show in Table 2, the integration of StaMo introduces negligible overhead, leaving the model’s inference speed almost entirely unaffected.
Table 3: Performance comparison of different methods (%).(*w/o wrist camera)
Method Spatial Object Goal Long Average
OpenVLA 80.2 81.3 75.8 49.7 71.7
OpenVLA + StaMo state 91.4 92.3 87.1 73.2 86
OpenVLA + StaMo motion 92.3 92.5 86.4 75.1 86.6
OpenVLA-OFT* 91.7 93.2 89.6 90.3 91.2
OpenVLA-OFT* + StaMo state 94.8 96.2 92.2 92.9 94.0
OpenVLA-OFT* + StaMo motion 92.3 94.1 90.2 92.1 92.1
@@
RDT (all Real) 91.6 93.3 86.7 73.3 86.2
RDT (1 Real) 71.7 70.0 66.7 43.3 62.9
RDT (1Real+ 4ATM) 83.3 81.7 71.7 56.7 73.4
RDT (1Real+ 4LAPA) 80.0 76.7 75.0 65.0 74.2
RDT (1Real + 4StaMo) 90.0 91.6 86.7 70.0 84.6
4.3 Policy Co-training Results
In Table 4, we detail the results of our policy co-training. Our approach utilizes a DIT-based model on the RDT architecture, trained on a per-task dataset of 10 robot trajectories and 40 video trajectories with StaMo-generated pseudo-actions. We evaluated the final model checkpoint for each task across 20 trials, averaging the success rate over three full runs. StaMo significantly outperforms both LAPA and ATM, suggesting that its pseudo-action labels are more faithful to the original actions.
4.4 Action Linear Probing Experiment
A central claim of our work is that an effective latent action can be formulated as the simple vector difference between two latent states encoded by StaMo. To quantitatively validate the quality and utility of these emergent latent actions, we employ a linear probing protocol (Alain & Bengio, 2016), a standard methodology for evaluating the efficacy of learned representations, to measure how much explicit, predictable information about the motion is contained within our latent action formulation.

### v2

# Mingyu Liu1,2,∗  Jiuhe Shu1,∗  Hui Chen1  Zeju Li1  Canyu Zhao1 Jiange Yang3  Shenyuan Gao4  Hao Chen1,†  Chunhua Shen1,5,†

@@ Abstract
A fundamental challenge in embodied intelligence is learning state representations that are both compact and expressive for world modeling and decision making, yet existing methods often remain either redundant or miss task-critical information. We propose an unsupervised approach that learns a highly compressed two-token state representation using a lightweight encoder and a pre-trained Diffusion Transformer (DiT) decoder, capitalizing on its strong generative prior. Our representation is efficient, interpretable, and integrates seamlessly into existing VLA-based models, improving performance by 11.6% on LIBERO and 31% in real-world task success rate with minimal inference overhead. More importantly, we find that the difference between these tokens, obtained via latent interpolation, naturally represents the motion, which can be further decoded into executable robot actions. This emergent capability reveals that our representation captures dynamics without explicit motion supervision. We name our method StaMo for its ability to learn generalizable robotic Motion from compact State representation, which is encoded from static images, challenging the prevalent dependence on learning robotic motions with complex temporal modeling and video data. Our learned representations also enhance policy co-training, outperforming prior methods by 10.4% with improved interpretability. Moreover, our approach scales effectively across diverse data sources, including real-world robot data, simulation, and human egocentric video. Project page at https://aim-uofa.github.io/StaMo/.
footnotetext: Equal contribution. Corresponding authors.
@@ 1 Introduction
Learning reusable and generalizable representations is a cornerstone of intelligent robotics systems. While visual representations in VLAs retain rich perceptual details for multimodal fusion, state representations for world modeling and intermediate reasoning serve a different purpose: they are designed to enable efficient future prediction and bridging visual planning with action execution. This role demands two key properties: First, it must be highly compact. Unlike rich visual embeddings used for perception, a state representation’s proximity to action generation demands a focus on actionable information to ensure efficient future prediction. Second, it must be expressive despite its compactness. Common low-dimensional approaches, such as trajectories [47] or flow fields [19, 50], often fail on this front; they capture basic motion but lack the semantic richness to encode goal states, interaction dynamics, or structured spatial relationships.
Building on these principles, we propose a compact state representation learned using a Diffusion Autoencoder fine-tuned on robotics data. Our method compresses an image into a highly compact latent state (as few as 2 tokens of 1024 dimensions). To ensure the representation is rich, we initialize its decoder from a powerful DiT model pretrained on internet data [18], reasoning that the ability to reconstruct pixels accurately requires an implicit understanding of key state information like robot poses and object interactions. This representation integrates seamlessly into existing VLM architectures, extensive experiments show that our method significantly enhances model performance with negligible overhead to inference speed, achieving improvements of + 11.6% in LIBERO and + 31% in real-world deployments.
Interestingly, from this compact state representation, we make an important discovery: robot motion (which is also known as latent action) naturally emerges from the state representation space. By simply performing a linear interpolation between the latent encodings of start and goal observations, we can generate smooth, plausible, and dynamically consistent motion trajectories. This observation provides an alternative to the dominant paradigm of learning latent actions from video data [53, 11, 52, 15]. While using video seems intuitive as action is inherently temporal, it presents significant challenges: it demands complex, computationally expensive temporal models and often produces ambiguous, coarse-grained actions. This is because high-variance motion within video clips causes models to learn averaged-out representations, making per-frame learning both inefficient and representationally flawed.
@@ 2 Related Works
2.1 Robot Representation Learning
Previous robotics research on visual representations has often faced a fundamental trade-off: methods excelling at motion representation, like latent actions [11, 53, 14, 6, 20], flow [19, 51], or trajectories [47], typically lack a rich, compact state representation. Conversely, approaches that encode detailed state information from raw images [46, 5] or dense features [55, 26, 34, 49, 32, 40] are often too high-dimensional and redundant to effectively represent motion via simple differences. StaMo, as shown in Figure 2, overcomes this dichotomy by achieving a unique balance. It learns to compress and encode a robot’s visual state into a highly compact token space that is both expressive enough for complex tasks and minimal enough that motion can be elegantly and powerfully represented as the difference between two states. This unified approach not only enables efficient world modeling but also demonstrates superior generalization and scalability, providing a robust foundation for future robotic systems.
2.2 World Modeling in VLA
Benefiting from powerful vision-language models (VLMs) [3, 16, 4] pre-trained on large-scale internet data, vision-language-action (VLA) models have demonstrated significant potential in generalizing across manipulation tasks. Typically, VLAs [8, 58, 13, 22, 25, 7, 48, 42, 45, 30, 43] map visual information and language instructions into the robot’s action space through end-to-end training. A natural extension is to endow the model with world modeling capabilities. Inspired by recent unified generative and comprehension models, a promising direction is to enable large models to reason with images—predicting future visual states while inferring actions or language, thereby creating a mutually reinforcing process. Several works [46, 55, 12, 27] have explored this idea. However, these approaches either require decoding full images during inference or rely on overly redundant state representations, limiting their generalization to novel scenes. In contrast, our method employs a more compact representation that improves model performance with minimal impact on inference speed. Furthermore, our representation exhibits strong generalization ability, enabling adaptation to unseen scenes without requiring further fine-tuning of the encoder.
@@ 2.3 Latent Action Learning
Robot learning scalability is constrained by scarce and heterogeneous robotic data, prompting the use of large-scale, action-free internet videos to instill foundational physical and operational knowledge for improved generalization and data efficiency [33]. Numerous studies have explored learning discrete [53, 15, 9, 11, 41, 10] or continuous [52, 20] latent actions from unlabeled data and have demonstrated their effectiveness across a wide range of downstream tasks. However, most of these approaches rely on carefully designed model architectures and depend on extracting frames from continuous videos, which introduces limitations such as sensitivity to frame sampling intervals, potential temporal biases, and poorly interpretable, blurry action representations. In contrast, our approach reveals that actionable representations can inherently emerge from large-scale natural image collections—challenging the conventional belief that latent actions must be learned from video sequences. We further demonstrate that such image-emergent actions exhibit superior interpretability and stronger generalization capabilities.
@@ 3 Method
3.1 Compress Image into Compact State
@@
To this end, as illustrated in Figure 1, the encoder of our Diffusion Autoencoder, denoted as $\mathcal{E}$, consists of a pretrained DINOv2 model [36] serving as a frozen feature extractor, which is followed by a transformer-based compressor, to map the input observations into a sequence of compact token representations. The decoder $\mathcal{D}$, is a Diffusion Transformer (DiT) [37] conditioned on these tokens to reconstruct the original image observations. Our implementation is built upon the Stable Diffusion 3, where only the transformer-based compressor and the DiT decoder are trained. The DINOv2 encoder remains frozen throughout the training process. We employ the same Flow Matching [2, 31, 28] objective used in base model training to optimize the model, which minimizes the loss between the velocity predicted by $\mathcal{D}$ and the target velocity $u$:
@@
where $\mathbf{x}_{0}$ is the input image observation and $\tau$ is the VAE encoder of pretrained diffusion model to produce latent $\mathbf{z}_{0}$. $\mathbf{z}_{t}=(1-\sigma_{t})\mathbf{z}_{0}+\sigma_{t}\epsilon$ is the linear interpolation between pure noise $\epsilon$ and $\mathbf{z}_{0}$. After training, The trained encoder $\mathcal{E}$ is able to yield the state representation $s$.
3.2 Interpolating Motion from States
@@ 3.3 StaMo for Efficient World Modeling
$L_{total}=\lambda_{action}L_{action}+\lambda_{future}(L_{mse}(s_{pred},s_{gt})\\ +L_{1}(s_{pred},s_{gt})),$ (2)
@@ 3.4 StaMo for Latent Motions Co-Training
We tested this co-training framework in Section 4.5, our method leads to substantial performance gains compared with previous method. The results confirm that the motions emerging from StaMo serve as an effective and robust proxy for true actions, providing a scalable pathway to enhance policy learning by leveraging the wealth of unlabeled visual data.
@@ 4.1 Qualitative Analysis of StaMo
StaMo demonstrates the ability to perform high-quality reconstructions of robotic manipulation images with strong generalization capabilities. It achieves excellent results on both in-domain and out-of-domain data. Quantitative reconstruction results are shown in Table 1, and some qualitative results are presented in Figure 3. More results can be found in Appendix. For main experiments, we use DROID [23] and LIBERO [29] as our training data.
@@ 4.2 World Modeling Results
In this section, we evaluate the application of our StaMo representation for efficient world modeling. A key advantage of our approach is its ability to integrate seamlessly into existing VLM-based frameworks. We demonstrate this by conducting experiments on two strong baselines: OpenVLA and its successor, OpenVLA-OFT. We specifically chose OpenVLA-OFT because it introduces architectural improvements like parallel decoding and action chunking, which result in a significantly longer action horizon. This requires the model to predict a state representation at a more distant future step, making it a more challenging and relevant benchmark for world modeling.
To validate our approach, we retrained both OpenVLA and OpenVLA-OFT, replacing their original vision encoders with our StaMo encoder and training them to predict StaMo states. We denote these modified architectures as OpenVLA* and OpenVLA-OFT*; the original, unmodified models are presented without the asterisk. For our OpenVLA-OFT* variant, we introduced a minor architectural modification, using only the third-person camera view and removing the wrist camera input. We evaluate the performance of our StaMo representation on the LIBERO benchmark to demonstrate its superior capability in assisting VLA manipulation through enhanced world modeling. For each task, we conduct 1,000 evaluation rollouts, and all results are presented as percentage success rates.
Table 2: Performance comparison of different methods (%).
Method Spatial Object Goal Long Average
DP [17] 78.3% 92.5% 68.3% 50.5% 72.4%
Octo-Base [42] 78.9% 85.7% 84.6% 51.1% 75.1%
Spatial-VLA [39] 88.2% 89.9% 78.6% 55.5% 78.1%
CoT-VLA [56] 87.5% 91.6% 87.6% 69.0% 83.9%
$\pi$0-FAST [38] 96.4% 96.8% 88.6% 60.2% 85.5%
WorldVLA [12] 87.6% 96.2% 83.4% 60.0% 81.8%
UniVLA [46] 95.4% 98.8% 93.6% 94.0% 96.5%
OpenVLA [25] 84.7% 88.4% 79.2% 53.7% 76.5%
OpenVLA* + DINOv2 Feature 88.6% 90.4% 83.5% 61.4% 80.9%
OpenVLA* + StaMo state 92.3% 94.8% 88.1% 75.2% 87.6%
OpenVLA* + StaMo motion 93.1% 95.1% 87.4% 76.9% 88.1%
OpenVLA-OFT [24] 93.7% 94.2% 89.7% 91.3% 92.2%
OpenVLA-OFT* + DINOv2 Feature 94.1% 95.0% 91.2% 93.0% 93.3%
OpenVLA-OFT* + StaMo state 96.8% 98.9% 95.0% 96.3% 96.8%
OpenVLA-OFT* + StaMo motion 95.3% 97.1% 94.1% 92.9% 94.9%
Both the StaMo state and motion representations can serve as predictive targets for a world model, and their performance is detailed in Table 2. Our findings indicate that the optimal choice of representation depends on the policy’s action horizon. For the standard OpenVLA, which predicts actions over a short, single-step interval, the motion representation yields better performance. This is because the latent motion is conceptually analogous to the delta end-effector (EE) pose used in frameworks like LIBERO, providing a direct and incremental target. Conversely, for OpenVLA-OFT, which plans over a longer execution step, the state representation is recommended, as it acts as a stable goal-conditioning signal.
4.3 Inference Speed Comparison
A key advantage of StaMo lies in its utilization of two highly compressed tokens to assist VLM-based models with world modeling. During the inference stage, it obviates the need to decode full images; the model only needs to predict the tokens themselves. This approach significantly enhances the efficiency of the world modeling process. We conducted a comparative analysis of the model’s inference frequency before and after integrating the StaMo representation, and also benchmarked its speed against other world modeling methods. Our findings indicate that StaMo introduces significantly less inference overhead. The comparison of different model’s inference frequency is presented in Table 3.
Table 3: Inference Speed Comparison.The table presents a comparison of inference speeds for different methods, with all models deployed on a single RTX 4090 GPU. We benchmark against other world model-based VLA approaches, such as UniVLA [46] and WorldVLA [12]. As the results show, our method achieves a higher inference frequency compared to previous methods, including directly using DINOv2 features.Our approach introduces minimal computational overhead compared to the original baseline model.
Method Frequency (Hz) $\uparrow$
WorldVLA [12] 2.27
UniVLA [46] 2.65
OpenVLA [25] 4.16
OpenVLA + DINOv2 Feature 2.86
@@ 4.3 Inference Speed Comparison
OpenVLA-OFT [24] 18.24
OpenVLA-OFT + DINOv2 Feature 11.42
OpenVLA-OFT + StaMo 17.82
Table 4: Zero-Shot Performance on Maniskill [21].
Method Pick Cube Push Cube Stack Cube Pull Cube Tool
OpenVLA [25] 28% 4% 18% 16%
UniVLA [46] 32% 7% 22% 20%
WorldVLA [12] 30% 10% 17% 22%
OpenVLA + StaMo state 41% 32% 48% 36%
4.4 Zero-Shot Transfer to New Environment
Due to the compact nature of the StaMo representation, which focuses on manipulation-relevant information rather than extraneous visual details, our model exhibits superior generalization capabilities. We validated this in Table 4 by taking the model trained on the LIBERO benchmark and applying it directly to the ManiSkill [21] environment. Crucially, all methods were deployed in this zero-shot transfer setting. We evaluated each task over 100 trials, with results presented as percentage success rates. The outcomes demonstrate that our approach achieves stronger performance than competing models under these conditions.
4.5 Policy Co-training Results
In Table 5, we detail the results of our policy co-training. Our approach utilizes a DIT-based model on the RDT architecture, trained on a per-task dataset of 10 robot trajectories and 40 video trajectories with StaMo-generated pseudo-actions. We evaluated the final model checkpoint for each task across 20 trials, averaging the success rate over three full runs. StaMo significantly outperforms both LAPA [53] and ATM [47], suggesting that its pseudo-action labels are more faithful to the original actions.
@@ 4.5 Policy Co-training Results
RDT (all Real) 91.6% 93.3% 86.7% 73.3% 86.2%
RDT (1 Real) 71.7% 70.0% 66.7% 43.3% 62.9%
RDT (1Real+ 4ATM [47]) 83.3% 81.7% 71.7% 56.7% 73.4%
RDT (1Real+ 4LAPA [53]) 80.0% 76.7% 75.0% 65.0% 74.2%
RDT (1Real + 4StaMo) 90.0% 91.6% 86.7% 70.0% 84.6%
4.6 Action Linear Probing Results
A central claim of our work is that an effective latent action can be formulated as the simple vector difference between two latent states encoded by StaMo. To quantitatively validate the quality and utility of these emergent latent actions, we employ a linear probing protocol [1], a standard methodology for evaluating the efficacy of learned representations, to measure how much explicit, predictable information about the motion is contained within our latent action formulation.